OpenStudy (hannahc234):

Solve the system of equations algebraically show all your steps y=x^2+2x y=3x+20

jhonyy9 (jhonyy9):

what is your opinion idea ?

OpenStudy (hannahc234):

I don't have one. I have no idea what to do here.

OpenStudy (hannahc234):

Could I use substitution?

jhonyy9 (jhonyy9):

yes right bc. the second case is y=3x+20 so you can substituting it in the first in place of y so try it in this way and see what will get

OpenStudy (hannahc234):

@jhonyy9 I've gotten to the step where I factor and solve for x and got X=-4 or X=10 I have no idea how to find corresponding y values using either originally equations could you please help me

OpenStudy (anonymous):

you can use substitution

jhonyy9 (jhonyy9):

3x+20=x^2+2x x^2 -x -20 =0 (x-5)(x+4)=0 hope this will help you sure right easy - yes ?

OpenStudy (hannahc234):

Yes

jhonyy9 (jhonyy9):

so than you know that x_1 = 5 and x_2 = -4 so than just you need substituting these value in place of x in case of one equation for example y=3x+20 and will get the y_1 and y_2 values right ok. ?

OpenStudy (hannahc234):

Okay

jhonyy9 (jhonyy9):

so than was my pleasure good luck bye bye

OpenStudy (hannahc234):

Thank you

OpenStudy (hannahc234):

How do you find the y1 and y2 values?

jhonyy9 (jhonyy9):

for y_1 substitute in place of x the x_1 and for y_2 the x_2 and that s all

OpenStudy (hannahc234):

Thank you so much

OpenStudy (hannahc234):

So the answer is (-4,8) and (5,35)?

jhonyy9 (jhonyy9):

yes right sure good job bye bye was my pleasure

OpenStudy (hannahc234):

Thanks so much you're the best